Circular No. 7913

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ig IN NGC 7424 E. M. Schlegel, Smithsonian Astrophysical Observatory and Chandra

X-ray Center; and S. Ryder, Anglo-Australian Observatory, report the detection of SN 2001ig (cf. IAUC 7772) with the Chandra X-ray Observatory. The 23400-s observation, obtained on 2002 May 21.98–22.27 UT, contains 30 counts, corresponding to an ACIS-S rate of 0.0013 count/s. For an adopted 3-keV thermal bremsstrahlung model absorbed by the Galactic column toward NGC 7424 (1.3 × 10^{20} cm^{−2}), the observed count rate corresponds to a 0.2–10.0-keV unabsorbed flux of 7 × 10^{−15} erg s^{−1} cm^{−2}. The unabsorbed flux changes by ∼ 20 percent for an adopted 1- or 5-keV temperature. Assuming a distance of 11 Mpc (Tully 1988, Nearby Galaxies Catalog, Cambridge University Press), the 0.2–10-keV luminosity is ∼ 1.2 × 10^{38} erg s^{−1}. An additional Chandra observation is scheduled to measure a light-curve trend. Continued ATCA monitoring (cf. IAUC 7777) indicates that, since peaking at a flux of ∼ 20 mJy at 4790 MHz in early Feb., SN 2001ig has remained almost unchanged, in contrast to the optically thin decline observed in most type-II radio supernovae.

COMETS C/2002 J8, C/2002 K5, C/2002 K6, C/2002 K7 (SOHO) Further to IAUC 7909, D. Hammer reports his measurements for addi-

tional Kreutz sungrazing comets found by K. Černis (C/2002 J8), M. Meyer (C/2002 K5), S. Hönig (C/2002 K6), and C. Dong (C/2002 K7) on SOHO website C2 images. C/2002 K6 was also visible in the C3 coronagraph. Reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2002-K72. The initial measurements of C/2002 J3 (noted on IAUC 7899) were by D. Biesecker.

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 J8 May 13.311  3^{h}21.^{m}1  +16^{o}48′
C/2002 K5  23.581  4 03.4  +19 03
C/2002 K6  24.321  4 06.2  +17 42
C/2002 K7  25.435  4 10.5  +19 23

CORRIGENDA On IAUC 7911, SN 2002cv, line 19, for dereddened read reddened On IAUC 7912, SN 2002dd, line 18, for 1′′.6 east read 1′′.6 west

2002 May 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green